2x-4y+2z=6
x-5y-5z=3
2x+y+4z=-2

Answer by Alan3354(69443)

2x-4y+2z=6
x-5y-5z=3
2x+y+4z=-2
-------------
You didn't ask a question or make a request.
---
If you want to solve for x, y & z:
I would multiply the 2nd eqn by 2, then eliminate x.
Then solve the 2 eqns for y & z.